# Hierarchical World Models as Visual Whole-Body Humanoid Controllers

### v2

@@ Abstract
Figure 1: Visual whole-body control for humanoids. We present Puppeteer, a hierarchical world model for whole-body humanoid control with visual observations. Our method produces natural and human-like motions without any reward design or skill primitives, and traverses challenging terrain.
@@ 1 Introduction
Due to the sheer complexity of such problems, prior work choose to make simplified assumptions, such as using low-dimensional (privileged) observations and actions (Heess et al., 2017; Peng et al., 2018; Wagener et al., 2022; Jiang et al., 2023), or (learned) skill primitives (Merel et al., 2017, 2018b; Hasenclever et al., 2020; Cheng et al., 2024). While these approaches are able to solve a variety of tasks to some extent, we expect to find a solution that is entirely data-driven and relies on as few assumptions as possible. In this paper, we propose to directly learn a controller on high-dimensional humanoid robot joints, leveraging model-based RL with an existing large-scale motion capture (MoCap) dataset (CMU, 2003).
@@
We propose a data-driven RL method for visual whole-body control that produces natural, human-like motions and can solve diverse tasks. Our approach, dubbed Puppeteer, is a hierarchical JEPA-style (LeCun, 2022) world model that consists of two distinct agents: a proprioceptive tracking agent that tracks a reference motion via joint-level control, and a visual puppeteer agent that learns to perform downstream tasks by synthesizing lower-dimensional reference motions for the tracking agent to track based on visual observations.
@@
Concretely, the two agents are trained independently in two separate stages using the model-based RL algorithm TD-MPC2 (Hansen et al., 2024) as a learning backbone. First, a single tracking world model is (pre)trained to track reference motions from pre-existing human MoCap data (CMU, 2003) re-targeted to a humanoid embodiment (Wagener et al., 2022). It learns a single model to convert any reference kinematic motion to physically executable actions. This is a departure from previous work that often learns multiple low-level models (Merel et al., 2017; Hasenclever et al., 2020). Importantly, this tracking agent can be saved and reused across all downstream tasks. In the second stage, we train a puppeteering world model that takes visual observation as inputs and outputs the reference motion for the tracking agent based on the specified downstream task. The puppeteer agent is trained with online environment interaction using the fixed tracking agent. A key feature of our framework is its striking simplicity: both world models are algorithmically identical (but differ in inputs/outputs) and can be trained using RL without any bells and whistles. Different from a traditional hierarchical RL setting, our puppeteer agent (high-level policy) outputs geometric locations for a small number of end-effector joints instead of a goal embedding. The tracking agent (low-level policy) is thus only required to learn joint-level physics. This makes the tracking agent easily sharable and generalizable across tasks, leading to an overall small computational footprint.
@@
To evaluate the efficacy of our approach, we propose a new task suite for visual whole-body humanoid control with a simulated 56-DoF humanoid, which contains a total of 8 challenging tasks. We show that our method produces highly performant control policies across all tasks compared to a set of strong model-free and model-based baselines: SAC (Haarnoja et al., 2018), DreamerV3 (Hafner et al., 2023), and TD-MPC2 (Hansen et al., 2024). Furthermore, we find that motions generated by our method are broadly preferred by humans in a user study with $46$ participants. We conclude the paper by carefully dissecting how each of our design choices influence results. Code for method and environments is available at https://nicklashansen.com/rlpuppeteer. Our main contributions can be summarized as follows:
@@
  • Task suite. We propose a new, challenging task suite for visual whole-body humanoid control with a simulated 56-DoF humanoid. The task suite has 8 tasks in total.
@@
  • Hierarchical world model. We propose a method for high-dimensional continuous control that leverages a learned hierarchical world model for planning.
@@
  • Evaluating “naturalness” of controllers. We develop several metrics for quantifying how natural and human-like generated motions are across tasks in our suite, including human preferences from a user study.
@@ 2 Preliminaries
TD-MPC2. We build upon the model-based reinforcement learning (MBRL) algorithm TD-MPC2 (Hansen et al., 2024), which represents the state-of-the-art in continuous control and has been shown to outperform alternatives in tasks with high-dimensional action spaces (Hansen et al., 2022, 2024; Sferrazza et al., 2024). Specifically, TD-MPC2 learns a latent decoder-free world model from environment interactions and selects actions by planning with the learned model. All components of the world model are learned end-to-end using a combination of joint-embedding prediction (Grill et al., 2020), reward prediction, and temporal difference (Sutton, 1998) losses, without decoding raw observations. During inference, TD-MPC2 follows the Model Predictive Control (MPC) framework for local trajectory optimization using Model Predictive Path Integral (MPPI) (Williams et al., 2015) as a derivative-free (sampling-based) optimizer. To accelerate planning, TD-MPC2 additionally learns a model-free policy prior which is used to warm-start the sampling procedure.
@@ 3 A Hierarchical World Model for High-Dimensional Control
A unique benefit of our approach is that a single tracking world model can be (pre)trained and reused across all downstream tasks. The tracking and puppeteering world models are algorithmically identical (but differ in inputs/outputs), and consist of the following 6 components:
@@
where $\mathbf{z}$ is a latent state. Because we consider episodic MDPs with termination conditions, we additionally add a termination prediction head $D$ (highlighted in Equation 1) that predicts the probability of termination conditioned on a latent state and action. Use of termination signals in the context of planning with a world model requires special care and has, to the best of our knowledge, not been explored in prior work; we introduce a novel method for this in Section 3.3. In the following, we describe the two agents and their interplay in the context of visual whole-body humanoid control. Minor implementation details are provided in Appendix B.
@@ 3.2 High-Level Puppeteering World Model
We now consider training a high-level puppeteering world model via online interaction in downstream tasks. As illustrated in Figure 2, the puppeteering model is trained (using downstream task rewards) to control the tracking model via commands $\mathbf{c}$, i.e., we redefine commands to now be the action space of the puppeteering agent. The tracking world model remains frozen (no weight updates) throughout this process, which allows us to reuse the same tracking model across all downstream tasks. Because the high-level agent uses planning for action selection, it natively supports temporal abstraction by outputting multiple commands $(\mathbf{c}_{t},\mathbf{c}_{t+1},\dots,\mathbf{c}_{t+H})$ for the low-level agent to execute; we treat the number of low-level steps per high-level step as a hyperparameter $k$ that allows us to trade strong motion prior (large $k$) for control granularity (small $k$).
@@ 3.3 Planning with Termination Conditions
Figure 4: Tasks. We develop 5 visual whole-body humanoid control tasks with a 56-DoF simulated humanoid (bottom), as well as 3 non-visual tasks (top). See Appendix C for more details.
@@ 4 Experiments
Our proposed method holds the promise of strong downstream task performance while still synthesizing natural and human-like motions. To evaluate the efficacy of our method, we propose a new task suite for whole-body humanoid control with multi-modal observations (vision and proprioceptive information) based on the “CMU Humanoid” model from DMControl (Tassa et al., 2018). Our simulated humanoid has 56 fully controllable joints ($\mathcal{A}\in\mathbb{R}^{56}$), and includes both head, hands, and feet. We aim to learn highly performant policies in a data-driven manner without the need for embodiment- or task-specific engineering (e.g., reward design, constraints, or auxiliary objectives), while synthesizing natural and human-like motions. Code for method and environments is available at https://nicklashansen.com/rlpuppeteer.
Figure 5: Learning curves. Episode return as a function of environment steps on all 8 tasks from our proposed task suite. Our method matches the performance of TD-MPC2 while producing more natural motions. We evaluate SAC and DreamerV3 only on proprioceptive tasks as they do not achieve any meaningful performance. Average of 3 random seeds; shaded area is $95\%$ CIs.
@@ 4.1 Experimental Details
Tasks. Our proposed task suite consists of 5 vision-conditioned whole-body locomotion tasks, and an additional 3 tasks without visual input. We provide an overview of tasks in Figure 4; they are designed with a high degree of randomization and include running along a corridor, jumping over hurdles and gaps, walking up the stairs, and circumnavigating obstacles (walls). All 5 visual control tasks use a reward function that is proportional to the linear forward velocity, while non-visual tasks reward displacement in any direction. Episodes are terminated at timeout ($500$ steps) or when a non-foot joint makes contact with the floor. We empirically observe that the TD-MPC2 baseline degenerates to highly unrealistic behavior without a contact-based termination condition.
Figure 6: Human preference in humanoid motions. Aggregate results from a user study ($n=46$) where humans are presented with pairs of motions generated by TD-MPC2 and our method, and are asked to provide their preference.
@@
Baselines. We benchmark our method against state-of-the-art RL algorithms for continuous control, including (1) widely used model-free RL method Soft Actor-Critic (SAC) (Haarnoja et al., 2018) which learns a stochastic policy and value function using a maximum entropy RL objective, (2) model-based RL method DreamerV3 (Hafner et al., 2020, 2021, 2023) which simultaneously learns a world model using a generative objective, and a model-free policy in the latent space of the learned world model, and (3) model-based RL method TD-MPC2 (Hansen et al., 2022, 2024) which learns a self-predictive (decoder-free) world model and selects actions by planning with the learned world model. Both our method and baselines use the same hyperparameters across all tasks. We provide further implementation details in Appendix B.
@@ 4.2 Main Results
We first present our main benchmark results, and then analyse and ablate each design choice.
Benchmark results. We evaluate our method, Puppeteer, and baselines on all 8 whole-body humanoid control tasks. Episode return as a function of environment steps is shown in Figure 5. We observe that the performance of our method is comparable to that of TD-MPC2 across all tasks (except stairs), whereas SAC and DreamerV3 does not achieve any meaningful performance within our computational budget of 3M environment steps. As we will soon reveal, TD-MPC2 produces better policies in terms of episode return on the stairs task, but far less natural behavior; this phenomenon is commonly referred to as reward hacking (Clark and Amodei, 2016; Skalse et al., 2022).
Table 1: Proxies for “naturalness”. Evaluated on the hurdles task. eplen denotes the average episode length over the course of training; height is the average torso height (gait) at end of training. Mean and std. across 3 seeds.
eplen $\uparrow$ height (cm) $\uparrow$
TD-MPC2 $70.7\pm 5.5$ $85.9\pm 4.7$
Ours $\mathbf{100.6\pm 1.0}$ $\mathbf{96\pm 0.2}$
@@ 4.3 Ablations & Analysis
Pretraining (tracking). Our method leverages both offline and online data during pretraining of the tracking world model. We ablate this training mixture in two distinct ways: (i) using only offline or online data, and (ii) reducing the number of unique MoCap clips seen during training. Interestingly, we find that leveraging both data sources leads to better tracking policies overall. We hypothesize that this is because offline data may help in learning to track especially difficult motions such as jumping and balancing on one leg, while online data improves state-action coverage and thus leads to a more robust world model overall. Similarly, training on more diverse MoCap clips also leads to better tracking performance. Training on all $836$ MoCap clips results in the best tracking world model, and we expect tracking to further improve with availability of more MoCap data.
Downstream tasks. We conduct three ablations that help us better understand the impact of a hierarchical approach to downstream tasks: (i) using a learned model-free policy in lieu of planning in either level of the hierarchy, (ii) pretraining of the high-level agent in addition to the low-level agent, and (iii) evaluating zero-shot generalization to unseen environment variations (gap length in the gaps task). The first two ablations are shown in Figure 8, and the latter is shown in Figure 9. We find that planning at both levels is critical to effective whole-body humanoid control, which we conjecture is due to the high dimensionality of the problem; this is supported by concurrent studies on high-dimensional continuous control with TD-MPC2 (Hansen et al., 2024; Sferrazza et al., 2024). Next, we pretrain agents on the corridor task and independently finetune on each visual control task. While the specific environments and motions differ between tasks, we find that our method benefits substantially from finetuning. We conjecture that this is because the need to control a low-level tracking agent is shared between all high-level agents, irrespective of the downstream task. Finally, we explore the zero-shot generalization ability of our method to harder, unseen variations of the gap task. Interestingly, we observe that our method generalizes to gap lengths up to 3×3$\times$3 × the training data without additional training. In light of these results, we believe that further investigation of the generalization ability of hierarchical world models will be a promising direction for future research.

### v3

@@ Abstract
Figure 1: Visual whole-body control for humanoids. We present Puppeteer, a hierarchical world model for humanoid control with visual observations. Our method produces natural and human-like motions without any reward design or skill primitives, and traverses challenging terrain.
@@ 1 Introduction
Due to the sheer complexity of such problems, prior work choose to make simplifying assumptions, such as using low-dimensional (privileged) observations and actions (Heess et al., 2017; Peng et al., 2018; Wagener et al., 2022; Jiang et al., 2023; Peng et al., 2021), or (learned) skill primitives (Merel et al., 2017; 2018b; Hasenclever et al., 2020; Peng et al., 2022; Cheng et al., 2024). Most related to our work, MoCapAct (Wagener et al., 2022) first learn $\sim$2600 individual tracking policies via RL, then distill them into a multi-clip tracking policy via imitation learning, and subsequently train a high-level RL policy to output goal embeddings for the multi-clip policy to track. While such approaches have been shown to transfer to simple reaching and velocity control tasks from proprioceptive inputs, we expect to find a solution that can perform complex, visual whole-body control tasks while remaining entirely data-driven and relying on as few assumptions as possible. In this paper, we propose to directly learn a visual controller for high-dimensional humanoid robot joints via model-based RL and an existing large-scale motion capture (MoCap) dataset (CMU, 2003), while requiring several orders of magnitude less interactions to learn new tasks compared to prior work.
@@
We propose a data-driven RL method for visual whole-body control that produces natural, human-like motions and can perform diverse tasks. Our approach, dubbed Puppeteer, is a hierarchical JEPA-style (LeCun, 2022) world model that consists of two distinct agents: a proprioceptive tracking agent that tracks a reference motion via joint-level control, and a visual puppeteer agent that learns to perform downstream tasks by synthesizing lower-dimensional reference motions for the tracking agent to track based on visual observations.
@@
Concretely, the two agents are trained independently in two separate stages using the model-based RL algorithm TD-MPC2 (Hansen et al., 2024) as a learning backbone. First, a single tracking world model is (pre)trained to track reference motions from pre-existing human MoCap data (CMU, 2003) re-targeted to a humanoid embodiment (Wagener et al., 2022). It learns a single model to convert any reference kinematic motion to physically executable actions. This is a departure from previous work that learns multiple low-level models (Merel et al., 2017; Hasenclever et al., 2020; Wagener et al., 2022). Importantly, this tracking agent can be saved and reused across all downstream tasks. In the second stage, we train a puppeteering world model that takes visual observation as inputs and outputs the reference motion for the tracking agent based on the specified downstream task. The puppeteer agent is trained with online environment interaction using the fixed tracking agent. A key feature of our framework is its striking simplicity: both world models are algorithmically identical (but differ in inputs/outputs) and can be trained using RL without any bells and whistles. Different from a traditional hierarchical RL setting, our puppeteer agent (high-level policy) outputs geometric locations for a small number of end-effector joints instead of a goal embedding. The tracking agent (low-level policy) is thus only required to learn joint-level physics. This makes the tracking agent easily sharable and generalizable across tasks, leading to an overall small computational footprint.
@@
To evaluate the efficacy of our approach, we propose a new task suite for visual whole-body humanoid control with a simulated 56-DoF humanoid, which contains a total of 8 challenging tasks. We show that our method produces highly performant control policies across all tasks compared to a set of strong model-free and model-based baselines: SAC (Haarnoja et al., 2018), DreamerV3 (Hafner et al., 2023), and TD-MPC2 (Hansen et al., 2024). Furthermore, we find that motions generated by our method are broadly preferred by humans in a user study with $51$ participants. We conclude the paper by carefully dissecting how each of our design choices influence results. Code for method and environments is available at https://www.nicklashansen.com/rlpuppeteer. Our main contributions can be summarized as follows:
@@
  • Task suite. We propose a new, challenging task suite for visual whole-body humanoid control with a simulated 56-DoF humanoid. The task suite has 8 tasks in total, and poses a significant challenge for existing state-of-the-art RL algorithms. At present, no such benchmark exists.
@@
  • Hierarchical world model. We propose a simple yet highly effective method for high-dimensional continuous control that uses a learned hierarchical world model for planning.
@@
  • Evaluating “naturalness” of controllers. We develop several metrics for quantifying how natural and human-like generated motions are across tasks in our suite, including human preferences from a user study. To the best of our knowledge, no prior work has explicitly evaluated naturalness of learned policies for humanoid control.
@@ 2 Preliminaries
TD-MPC2. We build upon the model-based reinforcement learning (MBRL) algorithm TD-MPC2 (Hansen et al., 2024), which represents the state-of-the-art in continuous control and has been shown to outperform alternatives in tasks with high-dimensional action spaces (Hansen et al., 2022; 2024; Sferrazza et al., 2024). Specifically, TD-MPC2 learns a latent decoder-free world model from environment interactions and selects actions by planning with the learned model. All components of the world model are learned end-to-end using a combination of joint-embedding prediction (Grill et al., 2020), reward prediction, and temporal difference (Sutton, 1998) losses, without decoding raw observations. During inference, TD-MPC2 follows the Model Predictive Control (MPC) framework for local trajectory optimization using Model Predictive Path Integral (MPPI) (Williams et al., 2015) as a derivative-free (sampling-based) optimizer. To accelerate planning, TD-MPC2 additionally learns a model-free policy prior which is used to warm-start the sampling procedure.
@@ 3 A Hierarchical World Model for High-Dimensional Control
A unique benefit of our approach is that a single tracking world model can be (pre)trained and reused across all downstream tasks. This is in contrast to much of prior work that either learn a large number (up to $\sim$2600) of low-level policies (Merel et al., 2017; 2018b; Hasenclever et al., 2020; Wagener et al., 2022), or train policies from scratch on each downstream task (Peng et al., 2018; 2021). The tracking and puppeteering world models are algorithmically identical (but differ in inputs/outputs), and consist of the following 6 components:
@@
where $\mathbf{z}$ is a latent state. Because we consider episodic MDPs with termination conditions, we additionally add a termination prediction head $D$ (highlighted in Equation 1) that predicts the probability of termination conditioned on a latent state and action. Use of termination signals in the context of planning with a world model requires special care and has, to the best of our knowledge, not been explored in prior work; we introduce a novel method for this in Section 3.3. In the following, we describe the two agents and their interplay in the context of visual whole-body humanoid control.
@@ 3.2 High-Level Puppeteering World Model
We now consider training a high-level puppeteering world model via online interaction in downstream tasks. As illustrated in Figure 2, the puppeteering model is trained (using downstream task rewards) to control the tracking model via commands $\mathbf{c}$, i.e., we redefine commands to now be the action space of the puppeteering agent. The tracking world model remains frozen (no weight updates) throughout this process, which allows us to reuse the same tracking model across all downstream tasks. Because the high-level agent uses planning for action selection, it natively supports temporal abstraction by outputting multiple commands $(\mathbf{c}_{t},\mathbf{c}_{t+1},\dots,\mathbf{c}_{t+H})$ for the low-level agent to execute; we treat the number of low-level steps per high-level step as a hyperparameter $k$ that allows us to trade strong motion prior (large $k$) for control granularity (small $k$). The high-level policy outputs commands at a fixed frequency regardless of whether the previous command was achieved.
@@ 3.3 Planning with Termination Conditions
Figure 4: Tasks. We develop 5 visual whole-body humanoid control tasks with a 56-DoF simulated humanoid (bottom), as well as 3 non-visual tasks (top). See Appendix E for more details.
@@ 4 Experiments
Our proposed method holds the promise of strong downstream task performance while still synthesizing natural and human-like motions. To evaluate the efficacy of our method, we propose a new task suite for whole-body humanoid control with multi-modal observations (vision and proprioceptive information) based on the “CMU Humanoid” model from DMControl (Tassa et al., 2018). Our simulated humanoid has 56 fully controllable joints ($\mathcal{A}\in\mathbb{R}^{56}$), and includes both head, hands, and feet. We aim to learn highly performant policies in a data-driven manner without the need for embodiment- or task-specific engineering (e.g., reward design, constraints, or auxiliary objectives), while synthesizing natural and human-like motions. Code for method and environments is available at https://www.nicklashansen.com/rlpuppeteer.
Figure 5: Learning curves. Episode return vs. environment steps on all 8 tasks from our proposed task suite. Our method generally matches the return of TD-MPC2 on these tasks while producing more natural motions. We only evaluate SAC and DreamerV3 on proprioceptive tasks as they do not achieve any meaningful performance. Average of 10 random seeds; shaded area is $95\%$ CIs.
@@ 4.1 Experimental Details
Tasks. Our proposed task suite consists of 5 vision-conditioned whole-body locomotion tasks, and an additional 3 tasks without visual input. We provide an overview of tasks in Figure 4; they are designed with a high degree of randomization and include running along a corridor, jumping over hurdles and gaps, walking up the stairs, and circumnavigating obstacles (walls). All 5 visual control tasks use a reward function that is proportional to the linear forward velocity, while non-visual tasks reward displacement in any direction. Episodes are terminated at timeout ($500$ steps) or when a non-foot joint makes contact with the floor. We empirically observe that the TD-MPC2 baseline degenerates to highly unrealistic behavior without a contact-based termination condition, and thus modify TD-MPC2 to support termination as described in Section 3.3. See Appendix E for details.
Figure 6: Human preference in humanoid motions. Aggregate results from a user study ($n=51$) where humans are presented with pairs of motions generated by TD-MPC2 and our method, and are asked to provide their preference.
@@
Baselines. We benchmark our method against state-of-the-art RL algorithms for continuous control, including (1) widely used model-free RL method Soft Actor-Critic (SAC) (Haarnoja et al., 2018) which learns a stochastic policy and value function using a maximum entropy RL objective, (2) model-based RL method DreamerV3 (Hafner et al., 2020; 2021; 2023) which simultaneously learns a world model using a generative objective, and a model-free policy in the latent space of the learned world model, (3) model-based RL method TD-MPC2 (Hansen et al., 2022; 2024) which learns a self-predictive (decoder-free) world model and selects actions by planning with the learned world model, (4) a hierarchical baseline that uses the same low-level TD-MPC2 agent as our method but trains a SAC policy as the high-level agent, and (5) the same hierarchical baseline but with DreamerV3 as the high-level agent (and TD-MPC2 at the low level). We refrain from making a direct comparison to MoCapAct (Wagener et al., 2022) and DeepMimic (Peng et al., 2018) as they do not support visual observations and require several orders of magnitude more environment interactions to learn downstream tasks. Both our method and baselines use the same hyperparameters across all tasks, as TD-MPC2 and DreamerV3 have been shown to be robust to hyperparameters across task suites (Hansen et al., 2024; Hafner et al., 2023; Sferrazza et al., 2024). For a fair comparison, we experiment with various design choices and hyperparameter configurations for SAC and report the best results that we obtained. We provide further implementation details in Appendix D.
@@ 4.2 Main Results
We first present our main benchmark results, and then analyze and ablate each design choice.
Benchmark results. We evaluate our method, Puppeteer, and baselines on all 8 whole-body humanoid control tasks. Episode return as a function of environment steps is shown in Figure 5. We observe that the performance of our method is comparable to that of TD-MPC2 across all tasks (except stairs), whereas SAC and DreamerV3 does not achieve any meaningful performance within our computational budget of 3M environment steps; hierarchical DreamerV3 achieves non-trivial yet still poor performance with TD-MPC2 as the low-level agent. As we will soon reveal, TD-MPC2 produces better policies in terms of episode return on the stairs task, but far less natural behavior (walking vs. rolling up stairs). We conjecture that this is a symptom of reward hacking (Clark & Amodei, 2016; Skalse et al., 2022). Sample videos are available at https://www.nicklashansen.com/rlpuppeteer.
Table 1: Proxies for “naturalness”. Evaluated on the gaps task. eplen denotes the average episode length (survival time) at 1M steps and at convergence; height is the average torso height (gait) at convergence. Mean and std. across 10 seeds.
eplen@1M $\uparrow$ eplen $\uparrow$ height (cm) $\uparrow$
TD-MPC2 $66.9\pm 9.8$ $181.6\pm 28.1$ $85.9\pm 4.7$
Ours $\mathbf{115.9\pm 5.2}$ $159.3\pm 5.9$ $\mathbf{96.0\pm 0.2}$
@@ 4.3 Ablations & Analysis
Pretraining (tracking). Our method leverages both offline and online data during pretraining of the tracking world model. We ablate this training mixture in two distinct ways: (i) using only offline or online data, and (ii) reducing the number of unique MoCap clips seen during training. Interestingly, we find that leveraging both data sources leads to better tracking policies overall. We hypothesize that this is because offline data may help in learning to track especially difficult motions such as jumping and balancing on one leg, while online data improves state-action coverage and thus leads to a more robust world model overall. Similarly, training on more diverse MoCap clips also leads to better tracking performance. Training on all $836$ MoCap clips results in the best tracking world model, and we expect tracking to further improve with availability of more MoCap data. See Appendix B for more results.
Downstream tasks. We conduct three ablations that help us better understand the impact of a hierarchical approach to downstream tasks: (i) using a learned model-free policy in lieu of planning in either level of the hierarchy, (ii) pretraining of the high-level agent in addition to the low-level agent, and (iii) evaluating zero-shot generalization to unseen environment variations (gap length in the gaps task). The first two ablations are shown in Figure 8, and the latter is shown in Figure 9. We find that planning at both levels is critical to effective whole-body humanoid control, which we conjecture is due to the high dimensionality of the problem; this is supported by concurrent studies on high-dimensional continuous control with TD-MPC2 (Hansen et al., 2024; Sferrazza et al., 2024). Next, we pretrain agents on the corridor task and independently finetune on each visual control task. While the specific environments and motions differ between tasks, we find that our method benefits substantially from finetuning. We conjecture that this is because the need to control a low-level tracking agent is shared between all high-level agents, irrespective of the downstream task. This is in contrast to contemporary work on pretraining that often pretrains on large-scale out-of-domain data (Hansen et al., 2023b; Xu et al., 2023). Finally, we explore the zero-shot generalization ability of our method to harder, unseen variations of the gap task. Interestingly, we observe that our method generalizes to gap lengths up to 3×3$\times$3 × the training data without additional training. In light of these results, we believe that further investigation of the generalization ability of hierarchical world models will be a promising direction for future research.